OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments)

OpenStudy (anonymous):

\[\frac{ 6a^3-13a^2+7a }{ 2a^3-5a+3a }\]

OpenStudy (anonymous):

this doesnt look solvable lol

hartnn (hartnn):

2 numbers with sum = -13 and product = 6*7 =42 are ?

OpenStudy (anonymous):

6 & 7?

hartnn (hartnn):

both negative :) -6, -7

OpenStudy (anonymous):

and -2 + -3 =5

OpenStudy (anonymous):

-5

hartnn (hartnn):

correct!

OpenStudy (anonymous):

so how do we solve this? lol

hartnn (hartnn):

factor the numerator \(a (6a^2-6a-7a+7) \\ a(...?...)\)

OpenStudy (anonymous):

a(6a-6-7...how would i factor the last 7, it doesn't have an a...

hartnn (hartnn):

factor 'a' only from first 2 terms factor -7 from last 2 terms

hartnn (hartnn):

factor 6a** from first 2 terms

OpenStudy (anonymous):

ooooh ok

hartnn (hartnn):

\(\large a(6a(a-1)-7(a-1))\) got that ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

factor out a-1 now

hartnn (hartnn):

\(\large a(6a(a-1)-7(a-1)) \\ \large a(a-1)(6a-7)\) got that ^ can you do the same thing for denominator?

OpenStudy (anonymous):

ill try lol

OpenStudy (anonymous):

a(2a^2-2a-3a+3)?

OpenStudy (anonymous):

a(2(a-1)-3(a-1))

OpenStudy (anonymous):

@ganeshie8